OpenStudy (anonymous):

5 moles of NaCl is how many grams

OpenStudy (anonymous):

5*40=200 gr

OpenStudy (anonymous):

Mw(NaCl) = 22.990 + 35.453 = 58.443 gram ie 1 mole weighs 58.443 g then 5 moles=58.443 * 5=292.215g